RC Meeting and Run Plan, Friday, January 10th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Target Readiness

  • Spin up measurements conducted this morning indicate masing. Upon inducing a gradient, the polarization increased. Need to do follow up tests in the transverse direction.
  • Do spin-up tests in the longitudinal direction overnight.
  • Target ladder alignment - target cell and reference cell look ok relative to the fiducial beam line (both upstream and downstream). For the carbon foils, the beam line is shifted down and to the right by 1.5 mm (while looking downstream).


Remaining work

In the enclosure:

  • Installing NMR legs and 2 BNC cables. -
  • Install EPR fiber bundle cable - later today
  • Install Reference cell RTDs - done
  • Check the long. + trans. spectral analyses + mounts - done
  • Cooling jets alignments - done
  • Reference cell gas connections - done
  • Install collimators - tomorrow
  • Follow the enclosure backout checklist
  • Check the angles on QWP in the laser enclosure.

Outside the enclosure:

  • Interlock wire broken - done
  • General clean up around the pivot area - tomorrow
  • Follow the hall backout checklist - tomorrow

Current Run Plan

  • Friday: Continue with spin up tests + follow remaining tasks listed above. Go into controlled access.
  • Saturday: Install collimators. Close up the enclosure. Start off with 5-pass and wait for MCC to checkout their equipment. Start taking 1-pass data for two days with the same settings as in December.
  • Sunday: 1-pass data
  • Monday: 1-pass data --> remove collimators and close the panels. Transition to 5-pass.

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]

Ongoing Work in the Hall

Planned Accelerator Configuration Changes

Special Requests

  • Keep target expert schedule filled out on whiteboard